The GNU time Command

Measuring Program Resource Usage Copyright \begin{matrix}c \\ ◯ \end{matrix} 1989, 1990 Free Software Foundation, Inc.

Permission is granted to make and distribute verbatim copies of this manual provided the copyright notice and this permission notice are preserved on all copies.

Permission is granted to copy and distribute modified versions of this manual under the conditions for verbatim copying, provided that the entire resulting derived work is distributed under the terms of a permission notice identical to this one.

Permission is granted to copy and distribute translations of this manual into another language, under the above conditions for modified versions, except that this permission notice may be stated in a translation approved by the Foundation.

1. time: Measuring Program Resource Usage

The time command runs another program, then displays information about the resources used by that program, collected by the system while the program was running. You can select which information is reported and the format in which it is shown (see ⟨undefined⟩ [Format], page ⟨undefined⟩), or have time save the information in a file instead of display it on the screen (see ⟨undefined⟩ [Usage], page ⟨undefined⟩).

1.1 Using the time command

The format of the time command is:

     time [-av] [-f format] [-o file] [+append] [+verbose]
     [+format format] [+output-file file] command [arg. . .]

time first runs the program command. When command finishes, time displays information about resources used by command, on the standard error output by default. If command exits with non-zero status, time displays a warning message and the exit status.

time determines which information to display about the resources used by the command from a format string (see ⟨undefined⟩ [Format], page ⟨undefined⟩). If no format is specified on the command line, but the TIME environment variable is set, its value is used as the format. Otherwise, a default format built into time is used (see ⟨undefined⟩ [Format], page ⟨undefined⟩).

Options to time must appear on the command line before command. Anything on the command line after command is passed as arguments to command.

‘-o file’

‘+output-file file’

Write the resource usage statistics to file instead of to the standard error stream. By default, this overwrites the file, destroying the file’s previous contents. This option is useful for collecting information on interactive programs and programs that produce output on the standard error stream.

‘-a’

‘+append’ This option is only useful with the ‘-o’ or ‘+output-file’ option; it causes time to append the information to the file instead of overwriting it.

‘-f format’ ‘+format format’

Use format as the format string that controls the output of time.

‘-v’
‘+verbose’

Use the built-in verbose format (different from the built-in default format), which displays every available piece of information on the program’s resource usage, on its own line with an English description of its meaning.

1.2 Formatting The Output

The format string controls the contents of the time output. The format string can be set using the ‘-f’ or ‘+format’ options, or a built-in verbose format can be selected using the ‘-v’ or ‘+verbose’ options; if they are not given, but the TIME environment variable is set, its value is used as the format string. Otherwise, a built-in default format is used. The default format is:

     %Uuser %Ssystem %Eelapsed %PCPU (%Xtext+%Ddata %Mmax)k
     %Iinputs+%Ooutputs (%Fmajor+%Rminor)pagefaults %Wswaps

The format string usually consists of resource specifiers interspersed with plain text. A percent sign (‘%’) in the format string causes the following character to be interpreted as a resource specifier, which is similar to the formatting characters in the C printf function.

A backslash (‘\’) introduces a backslash escape, which is translated into a single printing character upon output. ‘\t’ outputs a tab character, ‘\n’ outputs a newline, and ‘\\’ outputs a backslash. A backslash followed by any other character outputs a question mark (‘?’) followed by a backslash, to indicate that an invalid backslash escape was given.

Other text in the format string is copied verbatim to the output.

There are many resource specifications. Not all resources are measured by all versions of Unix, so some of the values might be reported as 0. Any character following a percent sign that is not listed in the table below causes a question mark (‘?’) to be output, followed by that character, to indicate that an invalid resource specifier was given.

The resource specification characters are:

‘%’ A literal ‘%’.

‘C’ The name and command line arguments of the command being timed. ‘D’ The average size of the process’s unshared data area, in Kilobytes. ‘E’ The elapsed real (wall clock) time used by the process, in [hours:]minutes:seconds.microseconds. ‘F’ Number of major, or I/O-requiring, page faults that occurred while the process was

running. These are faults where the page has actually migrated out of primary memory.

‘I’ Number of file system inputs by the process. ‘K’ The average total memory usage of the process, in Kilobytes. ‘M’ The maximum resident set size of the process during its lifetime, in Kilobytes. ‘O’ Number of file system outputs by the process. ‘P’ The percentage of the CPU that this job got. This is just user + system times divied

by the total running time.

‘R’ Number of minor, or recoverable, page faults. These are pages that are not valid (so they fault) but which have not yet been claimed by other virtual pages. Thus the data in the page is still valid but the system tables must be updated.

‘S’ Total number of CPU-seconds used by the system on behalf of the process (in kernel

mode), in seconds:microseconds.

‘U’ Total number of CPU-seconds that the process used directly (in user mode), in sec-

onds:microseconds.

‘W’ Number of times the process was swapped out of main memory. ‘X’ Average amount of shared text in the process, in Kilobytes. ‘Z’ The system’s page size, in bytes. This is a per-system constant, but varies between

systems.

‘c’ Number of times the process was context-switched involuntarily (because the time slice

expired).

‘k’ Number of signals delivered to the process. ‘p’ The average unshared stack size of the process, in Kilobytes. ‘r’ Number of socket messages received by the process. ‘s’ Number of socket messages sent by the process. ‘t’ The integral resident set size of the process, in Kilobytes. ‘w’ Number of times that the program was context-swapped voluntarily, for instance while

waiting for an I/O operation to complete.

‘x’ The exit status of the process.

The resource specification characters are a superset of those recognized by the tcsh builtin time command.

1.3 Examples

These examples assume that your command interpreter is bash. To run the command ‘wc /etc/hosts’ and show the default information:

     time wc /etc/hosts

To run the command ‘ls -Fs’ and show just the user, system, and total time:

     time -f "\t%R real,\t%U user,\t%S sys" ls -Fs

To edit the file bork and have time append the elapsed time and number of signals to the file ‘log’, reading the format string from the environment variable TIME:

     export TIME="\t%E,\t%k"
     time -a -o log emacs bork

1.4 Accuracy

The elapsed time is not collected atomically with the execution of the program; as a result, in bizarre circumstances (if the time command gets stopped or swapped out in between when the program being timed exits and when time calculates how long it took to run), it could be much larger than the actual execution time.

When the running time of a command is very nearly zero, some values (e.g., the percentage of CPU used) may be reported as either zero (which is wrong) or a question mark.

Most information shown by time is derived from the getrusage system call. The numbers are only as good as those returned by getrusage.

The ‘%I’ and ‘%O’ values are allegedly only “real” input and output and do not include those supplied by caching devices. The meaning of “real” I/O reported by ‘%I’ and ‘%O’ may be muddled for workstations, especially diskless ones.

Table of Contents

1 time: Measuring Program Resource Usage . . . . . . . . . . . . 1

1.1 Using the time command . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 1 1.2 Formatting The Output . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2 1.3 Examples. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .4 1.4 Accuracy . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 ii time: Measuring Program Resource Usage